OpenStudy (anonymous):

If g(x) = x2 − 3, find g(4).

OpenStudy (zzr0ck3r):

put a 4 where you see an x

OpenStudy (anonymous):

Dang it >:O

OpenStudy (anonymous):

curse you aqua scum!!!!!

OpenStudy (anonymous):

5? @zzr0ck3r

OpenStudy (anonymous):

where ever x is. you replace it with 4 and solve.

OpenStudy (zzr0ck3r):

so x^2-3 becomes 4^2-3

jimthompson5910 (jim_thompson5910):

I'm assuming x2 means x^2 x^2 means "x squared" so 4^2= 4 squared = 4*4 = 16

OpenStudy (anonymous):

5???????????????????????????

jimthompson5910 (jim_thompson5910):

and 4^2 is NOT equal to 8 you're thinking of 4*2 = 8

OpenStudy (anonymous):

what? now im confused @jim_thompson5910 @zzr0ck3r

jimthompson5910 (jim_thompson5910):

squaring something means you multiply the number by itself think of a square with side length of 4 

jimthompson5910 (jim_thompson5910):

other examples 5 squared = 5^2 = 5*5 = 25 8 squared = 8^2 = 8*8 = 64

OpenStudy (anonymous):

ohhh okay, now that makes sense !! @jim_thompson5910 thankyou !

jimthompson5910 (jim_thompson5910):

I'm glad it's clicking now

OpenStudy (anonymous):

Did you get your answer? G(4) = (4) ^2 - 3 G(4)=16-3 G(4) = 13